Simplify Answers 4 & 0 Why.
2y^2-8y=0

Answer by unlockmath(1688)

Hello,
This 2y^2-8y=0 can be rewritten as:
2y(y-4)=0 Now solve y to get:
2y=0 Divide both sides by 2 and we get:
y=0
and
y-4=0 Add 4 to both sides to get:
y=4
There we go!
Suggest you might get a book I wrote:
www.math-unlock.com